You plan to buy your "dream car" when you graduate in 12 months.
The cost of the car at that time will be $59,000 .
How much do you have to invest each month, starting next month, for 29 months to exactly pay for the car if your investments earn 5.00% APR (compounded monthly)?

Calculate your answer to the nearest dollar.

Write only the number with no dollar sign or comma (e.g., 3711)
Answer #1

Ans 1918

P = Periodic payments
r = rate of interest
n = no of years
Future Value of Annuity = P ( (1 + r)n - 1 ) / r
59000 = P* ((1 + 5%/12)^29 - 1) / (5%/12)
59000 = 30.7568586641773 * P
P = 59000 / 30.7568586641773
P = 1918